OpenStudy (anonymous):

Abby’s bus left at 2:00, going southbound at 55 mph. Bob’s went northbound starting at 3:00, going only 45 mph. How long will it take for them to get 420 miles apart? Round to the nearest tenth if necessary. 4.7 hours 4.2 hours 3.2 hours 3.6 hours

OpenStudy (anonymous):

Relative velocity is (55+45)=100mph Now try it from here on

OpenStudy (anonymous):

distance = velocity * time time = distance/velocity t = 420/100

OpenStudy (anonymous):

Solve the following for t: (55 t + 55) + 45 t = 420 t=73/20 or 3.62 hours

OpenStudy (anonymous):

Solve the following for t: (55 t + 55) + 45 t = 420 t=73/20 or 3.62 hours